Handheld fiber optic meters with white light polarization interferometry are useful for measuring temperature, pressure, and strain in electrically noisy environments. The fixed costs associated with manufacturing are $705,000 per year. If variable costs are $290 per unit and the company sells 4400 units per year, at what selling price per unit will the company break even?

The selling price at which the company will break even is determined to be $_____ per unit.
Answer #1

Break even price is that for which there are no profit or loss.
That is, total revenue = total cost.

Total revenue, TR = Price*Quantity sold = P*(4400) = 4400P
Total cost, TC = Fixed cost + Total variable cost = 705,000 + (290*4400) = 705,000 + 1,276,000 = 1,981,000

So, TR = TC gives,
4400P = 1,981,000
So, P = 1,981,000/4400
So, P = 450.23

The selling price at which the company will break even is determined to be $__450.23___ per unit.